# Words with factor complexity $2n+1$ and minimal critical exponent

### v2

@@ Abstract
Word $\mathbf{G}$} is the fixed point of the morphism $\gamma=[01,2,02]$. In 2019, Shallit and Shur showed that $\mathbf{G}$} has factor complexity $2n+1$. They also showed that $\mathbf{G}$} has critical exponent $\mu=2+\frac{1}{\lambda^{2}-1}=2.4808726\cdots$, where $\lambda=1.7548777$ is the real zero of $x^{3}-2x+x-1=0$. They conjectured that this was the least possible critical exponent among words with factor complexity $2n+1$. We confirm their conjecture. The proof, using an intricate case analysis, is by computer. The relevant program generates a ‘human readable’ proof.
@@ 1 Introduction
For several decades, the use of computer backtracking has been a standard tool in combinatorics on words. More recently, however, this area has seen a number of trends where computer-assisted proofs come closer to being ‘computer proofs’. One such trend has been seen in the continued success of the Walnut computer package [16, 22], which proves properties of automatic sequences. Over 100 papers using Walnut are listed by Shallit [23]. Of relevance to the current paper, in many situations Walnut can check the critical exponent of the fixed point of a morphism, as in a recent example by Baranwal et. al. [2].
@@
The critical exponent of a word is the supremum of the exponents of its factors. In recent years a large number of papers have been written answering questions of this form:
[…]
The various steps in the solution of Dejean’s Conjecture by several authors [12, 18, 17, 6, 3, 11, 20] answered this question for various classes of the form ${\cal L}=\Sigma_{k}^{\omega}$ where $\Sigma_{k}=\{0,1,2,\ldots,k-1\}$.
[…]
In this paper, we find the minimal critical exponent for the class of words with factor complexity $2n+1$. We confirm the 2019 conjecture of Shallit and Shur [24].
@@ 2 Word preliminaries
If $u$, $v\in\Sigma_{3}^{*}$, the concatenation $uv$ of $u$ and $v$ consists of the letters of $u$ followed by the letters of $v$. Thus if $u=012$ and $v=02$, we have $uv=01202$. We say that $u$ is a prefix of $uv$ and $v$ is a suffix of $uv$. We also say that $uv$ is a right extension of $u$ by $v$. If $u$, $v$, $w\in\Sigma_{3}^{*}$, we say that $v$ is a factor of uvwuvw. Thus $01$, $120$, and $02$ are respectively a prefix, factor, and suffix of $01202$, and $01202$ is a right extension of $01$ by $202$.
Suppose $u=u_{0}u_{1}u_{2}\cdots u_{n-1}$ where each $u_{i}\in\Sigma_{3}$. We say that positive integer $p$ is a period of $u$ if $u_{i}=u_{i+p}$ whenever $0\leq i\leq i+p\leq n-1$. In this case we say that $u$ is an $r$ power, where $r=n/p$. Thus 3 is a period of word $01201201$, which is a $8/3$ power. Call a word low if it contains no $r$ power for $r\geq 5/2$.
@@
Suppose $u=u_{0}u_{1}u_{2}\cdots u_{n-1}$ where each $u_{i}\in\Sigma_{3}$. The reverse of $u$ is the word $u^{R}=u_{n-1}\cdots u_{2}u_{1}u_{0}$. The reverse of morphism $h=[h(0),h(1),h(2)]$ is defined to be $h^{R}=[(h(0))^{R},(h(1))^{R},(h(2))^{R}]$. Thus, for example, $\gamma^{R}=[10,2,20]$. We see that for a word $u$ and morphism $h$ we have $(h(u))^{R}=h^{R}(u^{R}).$
As well as finite words, we consider right-infinite words over $\Sigma_{3}$. Formally, such a word is a sequence over $\Sigma_{3}$, ${\mathbf{u}}=\{u_{n}\}_{n=0}^{\infty}$ where the $u_{i}\in\Sigma_{3}$. We will use bold-face letters to distinguish right-infinite words from finite words. For each non-negative integer $n$, the length-$n$ prefix of $\mathbf{u}$} is the finite word $u_{0}u_{1}u_{2}\cdots u_{n-1}$. We say that $v\in\Sigma_{3}^{*}$ is a factor of $\mathbf{u}$} if it is a factor of some finite prefix of $\mathbf{u}$}. Call a right-infinite word low if each of its finite prefixes is low.
A final segment of a right-infinite word is the infinite analog of a suffix; a final segment of $\mathbf{u}$} is a right infinite word ${\mathbf{u}}^{\prime}=\{u_{n+i}\}_{n=0}^{\infty}$ for some non-negative integer $i$. We write ${\mathbf{u}}=u{\mathbf{u}}^{\prime}$ where is the prefix of $\mathbf{u}$} of length $i$.
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$. Fix a finite factor $b$ of $\mathbf{w}$}. Word $bu$ is a return word for $b$ in $\mathbf{w}$} if
  • Word bubbub is a factor of $\mathbf{w}$};
@@
Thus $bu$ is a return word for $b$ in $\mathbf{w}$} if bubbub is a shortest word ‘getting back to’ $b$. If $b$ is a factor of $\mathbf{w}$} then a final segment of $\mathbf{w}$} can be concatenated from the return words of $b$.
The set of all right-infinite words over $\Sigma_{3}$ is denoted by $\Sigma_{3}^{\omega}$. Right infinite words arise in various ways: If $u=u_{0}u_{1}\cdots u_{p-1}$ where $p$ is positive and the $u_{i}$ are letters, then $u^{\infty}=\{u_{i}\}_{i=0}^{\infty}$ where $u_{i}=u_{i\%p}$ where $i\%p$ is the remainder after $i$ is divided by $p$. Thus, for example,
@@
Another way a right-infinite word arises is via certain morphisms. Let $h$ be a morphism such that $a$ is a prefix of $h(a)$ for some $a\in\Sigma_{3}$. By induction, $h^{n-1}(a)$ is a prefix of $h^{n}(a)$ for each $n$. If $|h^{n}(a)|$ increases without bound, we can define $h^{\omega}(a)$ to be the unique right-infinite word having $h^{n}(a)$ as a prefix for each $n$. We refer to $h^{\omega}(a)$ as a fixed point of $h$.
In this paper we will be particularly interested in the fixed point ${\mathbf{G}}=\gamma^{\omega}(0)$.
A third (non-constructive) way to obtain right-infinite words is via König’s Tree Lemma: A subset $L\subseteq\Sigma_{3}^{*}$ is said to be a factorial language if whenever $\ell$ is in $L$, every factor of $\ell$ is also in $L$. Here is one formulation of König’s Tree Lemma in the combinatorics on words context:
Lemma 1 (König’s Tree Lemma).
@@
We say that right-infinite word ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ has letter frequencies $\rho_{0}$, $\rho_{1}$, $\rho_{2}$, if for any $\epsilon_{1}>0$ there is an $N_{1}$ such that if $u$ is a factor $\mathbf{u}$} of length at least $N_{1}$, then
@@
Using standard methods, it can be shown that $\mathbf{G}$} has letter frequencies. (For values of the $\rho_{a}$ see the entry for sequence A287104 in the Online Encyclopedia of Integer Sequences; the word $\mathbf{G}$} is a recoding of this sequence.)
@@ 3 Main Theorem
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ have factor complexity $2n+1$. Then the critical exponent of $\mathbf{w}$} is at least $\mu$.
@@
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ be low, and let $\mathbf{w}$} have factor complexity $2n+1$. There is a non-erasing morphism $h$ such that either:
  • For every factor $g$ of $\mathbf{G}$}, $h(g)$ is a factor of $\mathbf{w}$}.
  • For every factor $g$ of $\mathbf{G}$}, $h(g^{R})$ is a factor of $\mathbf{w}$}.
The Main Theorem will then follow from this lemma:
@@
Let ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ have irrational critical exponent $\nu$, $2<\nu<3.$ Suppose that $\mathbf{u}$} has letter frequencies $\rho_{0}$, $\rho_{1}$, $\rho_{2}$, such that for any $\epsilon_{1}>0$ there is an $N_{1}$ such that if $u$ is a factor $\mathbf{u}$} of length at least $N_{1}$, then
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ and let $h:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ be a non-erasing morphism. Suppose that for each factor $g$ of $\mathbf{u}$}, either $h(g)$ or $h(g^{R})$ is a factor of $\mathbf{w}$}. Then the critical exponent of $\mathbf{w}$} is at least $\nu$.
@@
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ have factor complexity $2n+1$. Note that $\mu<5/2$. Thus if $\mathbf{w}$} is not low, then the critical exponent of $\mathbf{w}$} is at least $5/2>\mu$, and we are done.
Suppose then that $\mathbf{w}$} is low. By the Structure Theorem, there is a non-erasing morphism $h$ such that for every factor $g$ of $\mathbf{G}$}, either $h(g)$ is a factor of $\mathbf{w}$} or $h(g^{R})$ is a factor of $\mathbf{w}$}. (This is actually weaker than the conclusion of the Structure Theorem).
Word $\mathbf{G}$} has letter frequencies.
It is easy to show that $\mu$ is irrational.
By Lemma 2, then, the critical exponent of ${\mathbf{w}})$ is at least $\mu$. ∎
@@ 4 Eligible words and their factors
Suppose ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is eligible. Then every final segment of $\mathbf{w}$} contains letters 0, 1, and 2.
Proof.
It suffices to show that every final segment contains letter 0.
Suppose some final segment contains only letters 1 and 2. It must contain factors 12 and 21; otherwise it would end with $1^{\omega}$ or $2^{\omega}$ and therefore contain $5/2$ powers.
It must also contain factor 11 (and, symmetrically, factor 22): If not, assume it starts with 1, replacing it with one of its final segments if necessary. Then it is concatenated from the return words of 1. Since 222 is a $3$ power, the only possible return words of 1 in $\mathbf{w}$} are 12 and 122. The return word 12 can only be used once in this concatenation, or else it appears in the context $2\underline{12}12$, which is a $5/2$ power. Then a final segment is $(122)^{\omega}$, which contains $5/2$ powers. This is a contradiction. Thus 11 must indeed occur.
So far we have shown that every final segment of $\mathbf{w}$} contains factors 11, 12, 21, and 22. Since $\mathbf{w}$} contains $2(2)+1=5$ factors of length 2, there will be exactly one length-2 factor of $\mathbf{w}$} containing a 0. It follows that the only 0 in $\mathbf{w}$} is its first letter. We thus see that $\mathbf{w}$} contains no factor of the form $a0$ where $a\in\Sigma_{3}$.
Consider the following set of 10 length-4 words:
$S=\{1121,1122,1211,1212,1221,2112,2121,2122,2211,2212\}.$
Let $s$ and $t$ be distinct words of $S$. Suppose $u$ is a low word over $\Sigma_{3}$ such that the factor set of $u$ does not include any of 00, 10, 20, $s$, or $t$. For each of the $\binom{10}{2}$ possibilities for $s$ and $t$, a backtrack search shows that the length of $u$ is at most 39. It follows then that $\mathbf{w}$} contains 9 out of 10 words of $S$ as factors. In addition, $\mathbf{w}$} contains a length-4 factor starting with 0. However, $\mathbf{w}$} is eligible, and cannot contain $10$ length-4 factors. This is a contradiction. ∎
Suppose that $\mathbf{w}$} is eligible. Consider the directed graph $D$ with vertex set $\Sigma_{3}$, and with a directed edge $xy$ for $x,y\in\Sigma_{3}$ exactly when $xy$ is a factor of $\mathbf{w}$}. Thus $\mathbf{w}$} can be walked on $D$. Since $\mathbf{w}$} has exactly 5 length-2 factors, $D$ has exactly 5 edges.
By the previous lemma, graph $D$ must be strongly connected. Therefore, it must have at least 3 non-loop edges, and if it has exactly three non-loop edges, they form a directed cycle.
It is convenient to keep track of non-factors of $\mathbf{w}$}: Let $F=\{u\in\Sigma_{3}^{*}:u\text{ is not a factor of }{\mathbf{w}}\}.$ To prove the Structure Theorem, it suffices to consider two cases involving $F$:
@@
The following lemmas give sufficient conditions for concluding that the Structure Theorem holds:
@@
Suppose that ${\mathbf{w}}\in\{0,1,2\}^{\omega}$ is low. Suppose that $\mathbf{w}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\{0,1,2\}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
@@
    Word $b$ is a prefix of $a$, which is a prefix of $c$.
@@
    Words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
Then every factor of $h({\mathbf{G}})$ is a factor of $\mathbf{w}$}.
@@
Suppose that ${\mathbf{z}}\in\{0,1,2\}^{\omega}$ is low. Suppose that a final segment of $\mathbf{z}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\{0,1,2\}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
@@
    Word $b$ is a suffix of $a$, which is a suffix of $c$.
@@
    Words $b$ and $c$ have a common prefix $p$ such that $|p|\geq|b|/2$.
Then if $g$ is any factor of $\mathbf{G}$}, then $h(g^{R})$ is a factor of $\mathbf{z}$}.

### v3

@@ Abstract
Let the word $\mathbf{G}$} be the fixed point of the morphism $\gamma$ sending 0 to $01$, 11 to $2$, and $2$ to $02$. In 2019, Shallit and Shur showed that $\mathbf{G}$} has factor complexity $2n+1$. They also showed that $\mathbf{G}$} has critical exponent $\mu=2+\frac{1}{\lambda^{2}-1}=2.4808726\cdots$, where $\lambda=1.7548777$ is the real zero of $x^{3}-2x+x-1=0$. They conjectured that this was the least possible critical exponent among the words with factor complexity $2n+1$. We confirm their conjecture. We anticipate that our method, including an intricate case analysis by computer, will have wider application.
@@ 1 Introduction
The critical exponent of a word is the supremum of the exponents of its factors. (Definitions of basic terms in combinatorics on words, such as factors, exponents, etc., follow in the next section.) In recent years a large number of papers have been written answering questions of the following form:
[…]
For example, the minimum critical exponent for a Sturmian word is
$2+\varphi=3.618033988749\cdots$
where $\varphi$ is the golden ratio. This is realized by the Fibonacci word, the fixed point of the morphism $\Phi$ sending 0 to $01$ and 11 to 0. One starting point for the study of Sturmian words and their generalizations is the relevant chapter in Allouche and Shallit’s book [1].
Words of factor complexity $2n+1$ are considered in the paper of Cassaigne, Labbé, and Leroy [6]. Such words are connected to the well-studied Arnoux-Rauzy words: the Arnoux-Rauzy words over a $k$-letter alphabet have factor complexity (k1)n+1(k-1)n+1. Note that the Arnoux-Rauzy words over a binary alphabet are just the Sturmian words. In the present paper, we find the minimal critical exponent for the class of words with factor complexity $2n+1$. This confirms the 2019 conjecture of Shallit and Shur [28]. Taken together with Carpi and De Luca’s result [5] for Sturmian words, this gives the beginnings of a Dejean-type theorem for Arnoux-Rauzy words: we know the minimal critical exponent in the cases $k=2$ and $k=3$.
To find the minimal critical exponent for binary rich words, it was necessary to prove the following structure theorem:
Theorem 1.
[9] Let ${\mathbf{w}}\in\Sigma_{2}^{\omega}$ be a $14/5$-power-free rich word. For every $n\geq 1$, a suffix of $\mathbf{w}$} has the form $f(\phi^{n}(w_{n}))$ or $f(g(\phi^{n}(w_{n})))$ for some word $w_{n}\in\Sigma_{3}^{\omega}$ where
$f(0)=$ $0$ $g(0)=$ $011$ $\phi(0)=$ $01$
$f(1)=$ $01$ $g(1)=$ $0121$ $\phi(1)=$ $02$
$f(2)=$ $011$ $g(2)=$ $012121$ $\phi(2)=$ $022.$
Paraphrasing, if we let $\mathbf{H}$} be the fixed point of $\phi$, then the infinite $14/5$-power-free rich binary words look like one of $h_{1}({\mathbf{H}})$ and $h_{2}({\mathbf{H}})$, where $h_{1}=f$, and $h_{2}=f\circ g$.
In the present paper, we prove the analogous structure theorem for the infinite words with factor complexity $2n+1$, and not containing $r$-powers for $r\geq 5/2$. However, instead of the two possible morphisms $h_{1}$ and $h_{2}$ in the above-mentioned theorem, there are several possible morphisms. (See Figures 1 and 2 and following.) The case analysis is therefore complex. (Interestingly, Shallit and Shur [28] prove that the minimum critical exponent for the infinite words with factor complexity $2n$ is exactly $5/2$.)
For several decades, the use of computer backtracking has been a standard tool in combinatorics on words. More recently, however, this area has seen a number of trends where computer-assisted proofs come closer to being ‘computer proofs’. One such trend has been seen in the continued success of the Walnut computer package [20, 26], which proves properties of automatic sequences. Over 100 papers using Walnut are listed by Shallit [27]. Of relevance to the current paper, in many situations Walnut can check the critical exponent of the fixed point of a morphism, as in a recent example by Baranwal et al. [3].
@@ 2 Word preliminaries
If $u$, $v\in\Sigma_{3}^{*}$, the concatenation $uv$ of $u$ and $v$ consists of the letters of $u$ followed by the letters of $v$. Thus if $u=012$ and $v=02$, we have $uv=01202$. We say that $u$ is a prefix of $uv$ and $v$ is a suffix of $uv$. We also say that $uv$ is a right extension of $u$ by $v$. If $u$, $v$, $w\in\Sigma_{3}^{*}$, we say that $v$ is a factor of uvwuvw. If word $y$ is not a factor of word $z$ we say that $z$ omits the factor $y$. Thus $01$, $120$, and $02$ are respectively a prefix, factor, and suffix of $01202$, and $01202$ is a right extension of $01$ by $202$. Word $01202$ omits the factor $10$.
Suppose $u=u_{0}u_{1}u_{2}\cdots u_{n-1}$ where each $u_{i}\in\Sigma_{3}$. We say that positive integer $p$ is a period of $u$ if $u_{i}=u_{i+p}$ whenever $0\leq i\leq i+p\leq n-1$. In this case we say that $u$ is an $r$ power, where $r=n/p$. Thus 3 is a period of the word $01201201$, which is an $8/3$ power. Call a word low if it contains no $r$ power for $r\geq 5/2$.
@@
Suppose $u=u_{0}u_{1}u_{2}\cdots u_{n-1}$ where each $u_{i}\in\Sigma_{3}$. The reverse of $u$ is the word $u^{R}=u_{n-1}\cdots u_{2}u_{1}u_{0}$. The reverse of the morphism $h=[h(0),h(1),h(2)]$ is defined to be $h^{R}=[(h(0))^{R},(h(1))^{R},(h(2))^{R}]$. Thus, for example, $\gamma^{R}=[10,2,20]$. We see that for a word $u$ and the morphism $h$ we have $(h(u))^{R}=h^{R}(u^{R}).$
As well as the finite words, we consider the right-infinite words over $\Sigma_{3}$. Formally, such a word is a sequence ${\mathbf{u}}=\{u_{n}\}_{n=0}^{\infty}$ where each $u_{i}$ is a single letter. We use bold-face letters to distinguish right-infinite words from finite words. For each non-negative integer $n$, the length-$n$ prefix of $\mathbf{u}$} is the finite word $u_{0}u_{1}u_{2}\cdots u_{n-1}$. We say that $v\in\Sigma_{3}^{*}$ is a factor of $\mathbf{u}$} if it is a factor of some finite prefix of $\mathbf{u}$}. Call a right-infinite word low if each of its finite prefixes is low.
We extend the notaion of suffix to right-infinite words; a suffix of $\mathbf{u}$} is a right-infinite word ${\mathbf{u}}^{\prime}=\{u_{n+i}\}_{n=0}^{\infty}$ for some non-negative integer $i$. We write ${\mathbf{u}}=u{\mathbf{u}}^{\prime}$ where $u$ is the prefix of $\mathbf{u}$} of length $i$.
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$. Fix a finite factor $b$ of $\mathbf{w}$}. The word $bu$ is a $b$-block of $\mathbf{w}$} if
  • The word bubbub is a factor of $\mathbf{w}$};
@@
If $b$ appears infinitely often as a factor of $\mathbf{w}$} then a suffix of $\mathbf{w}$} can be concatenated from the $b$-blocks of $\mathbf{w}$}.
The set of all right-infinite words over $\Sigma_{3}$ is denoted by $\Sigma_{3}^{\omega}$. Right-infinite words arise in various ways: If $u=u_{0}u_{1}\cdots u_{p-1}$ where $p$ is positive and the $u_{i}$ are letters, then $u^{\infty}=\{u_{i}\}_{i=0}^{\infty}$ where $u_{i}=u_{i\bmod p}$. Thus, for example,
@@
Another way a right-infinite word arises is via certain morphisms. Let $h$ be a morphism such that $a$ is a prefix of $h(a)$ for some $a\in\Sigma_{3}$. By induction, $h^{n-1}(a)$ is a prefix of $h^{n}(a)$ for each $n$. If $|h^{n}(a)|$ increases without bound, we can define $h^{\omega}(a)$ to be the unique right-infinite word having $h^{n}(a)$ as a prefix for each $n$. For a right-infinite word $\mathbf{u}$} over $\Sigma_{3}$, we define $h({\mathbf{u}})$ to be the right-infinite word having $h(u)$ as a prefix for each prefix $u$ of $\mathbf{u}$}; thus $h^{\omega}(a)$ is a fixed point of $h$. In this paper we are particularly interested in the fixed point ${\mathbf{G}}=\gamma^{\omega}(0)$.
A third (non-constructive) way to obtain right-infinite words is via König’s lemma [18]. Here is one formulation of König’s lemma in the combinatorics on words context: A subset $L\subseteq\Sigma_{3}^{*}$ is said to be a factorial language if whenever $\ell$ is in $L$, every factor of $\ell$ is also in $L$.
Lemma 1 (König’s lemma).
@@
We say that the right-infinite word ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ has letter frequencies $\rho_{0}$, $\rho_{1}$, $\rho_{2}$, if for all $\epsilon_{1}>0$ there is an $N_{1}$ such that if $u$ is a factor $\mathbf{u}$} of length at least $N_{1}$, then
@@
Using standard methods, it can be shown that $\mathbf{G}$} has letter frequencies. (For values of the $\rho_{a}$ see the entry for sequence A287104 in the On-Line Encyclopedia of Integer Sequences; the word $\mathbf{G}$} is a recoding of this sequence.)
@@ 3 Main Theorem
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ have factor complexity $2n+1$. Then the critical exponent of $\mathbf{w}$} is at least $\mu=2+\frac{1}{\lambda^{2}-1}=2.4808726\cdots$, where $\lambda=1.7548777$ is the real zero of $x^{3}-2x+x-1=0$.
@@
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ be eligible. There is a non-erasing morphism $h$ such that either
  • For every factor $g$ of $\mathbf{G}$}, the word $h(g)$ is a factor of $\mathbf{w}$}.
  • For every factor $g$ of $\mathbf{G}$}, the word $h(g^{R})$ is a factor of $\mathbf{w}$}.
The main theorem follows from this lemma:
@@
Let ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ have irrational critical exponent $\nu$, with $2<\nu<3.$ Suppose that $\mathbf{u}$} has letter frequencies $\rho_{0}$, $\rho_{1}$, $\rho_{2}$. Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ and let $h:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ be a non-erasing morphism. Suppose that for each factor $g$ of $\mathbf{u}$}, either $h(g)$ or $h(g^{R})$ is a factor of $\mathbf{w}$}. Then the critical exponent of $\mathbf{w}$} is at least $\nu$.
@@
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ have factor complexity $2n+1$. Recall that $\mu=2+\frac{1}{\lambda^{2}-1}=2.4808726\cdots$, where $\lambda=1.7548777$ is the real zero of $x^{3}-2x+x-1=0$. Thus $\mu<5/2$. If $\mathbf{w}$} is not low, then the critical exponent of $\mathbf{w}$} is at least $5/2>\mu$, and we are done.
Suppose then that $\mathbf{w}$} is low. By the structure theorem, there is a non-erasing morphism $h$ such that for every factor $g$ of $\mathbf{G}$}, either $h(g)$ is a factor of $\mathbf{w}$} or $h(g^{R})$ is a factor of $\mathbf{w}$}. (This is actually weaker than the conclusion of the structure theorem). The word $\mathbf{G}$} has letter frequencies. It is easy to show that $\mu$ is irrational. By Lemma 2, then, the critical exponent of $\mathbf{w}$} is at least $\mu$. ∎
@@ 4 Eligible words and their factors
Suppose ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is eligible. Then every suffix of $\mathbf{w}$} contains letters 0, 1, and 2.
Suppose that $\mathbf{w}$} is eligible. Consider the directed graph $D$ with vertex set $\Sigma_{3}$, and with a directed edge $xy$ for $x,y\in\Sigma_{3}$ exactly when $xy$ is a factor of $\mathbf{w}$}. Then there is a directed path on $D$ labeled with the letters of $\mathbf{w}$}. Since $\mathbf{w}$} has exactly 5 length-2 factors, the graph $D$ has exactly 5 edges.
By the previous lemma, the graph $D$ must be strongly connected; that is, there is a directed path between any two vertices of $D$. Therefore, $D$ must have at least 3 non-loop edges, and if it has exactly three non-loop edges, they form a directed cycle.
It is convenient to keep track of words which are not factors of $\mathbf{w}$}: Let $F=\{u\in\Sigma_{3}^{*}:u\text{ is not a factor of }{\mathbf{w}}\}.$ To prove the structure theorem, it suffices to consider two cases involving $F$.
@@
The following technical lemmas give sufficient conditions for concluding that the structure theorem holds:
@@
Suppose that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is low. Suppose that some suffix of $\mathbf{w}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that
@@
    The word $b$ is a prefix of $a$, which is a prefix of $c$.
@@
    The words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
If $g$ is any factor of $\mathbf{G}$}, then $h(g)$ is a factor of $\mathbf{w}$}.
@@
Suppose that ${\mathbf{z}}\in\Sigma_{3}^{\omega}$ is low. Suppose that a suffix of $\mathbf{z}$} has the form $h({\mathbf{u}})$ for some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$, where $h=[a,b,c]$ for some $a,b,c\in\Sigma_{3}^{+}$ such that
@@
    The word $b$ is a suffix of $a$, which is a suffix of $c$.
@@
    The words $b$ and $c$ have a common prefix $p$ such that $|p|\geq|b|/2$.
If $g$ is any factor of $\mathbf{G}$}, then $h(g^{R})$ is a factor of $\mathbf{z}$}.
Lemma 6 (First Parsing Lemma).
Let $S\subseteq\Sigma_{3}^{*}$ and let $b\in\Sigma_{3}^{*}$. Let $d$ be a non-negative integer. Suppose that every low word of length $d+1$ contains a factor from $S\cup\{b\}$. Then if $\mathbf{w}$} is an eligible right-infinite word which omits the words of $S$, a suffix of $\mathbf{w}$} can be concatenated from $b$-blocks of the form $bu$, where $|u|\leq d$.
Lemma 7 (Second Parsing Lemma).
Let $S,B\subseteq\Sigma_{3}^{*}$. Let $\mathbf{w}$} be an eligible right-infinite word which omits the words of $S$, such that $\mathbf{w}$} is concatenated from words of $B$. Let $s\in B$ have the property that for all $r,t\in B$, either rstrst is not low, or rstrst contains a factor from $S$. Then a suffix of $\mathbf{w}$} is concatenated from words of $B-\{s\}$.